# Global Compression Commander: Plug-and-Play Inference Acceleration for High-Resolution Large Vision-Language Models

### v4

@@ 5.4 Efficiency Analysis
Table 4 compares inference efficiency among SparseVLM, FasterVLM and our GlobalCom2. Unlike SparseVLM which requires explicit attention scores during decoding and is incompatible with FlashAttention [10], FasterVLM and our GlobalCom2 enable efficient computation before LLM decoding for efficient computation. As a plug-and-play solution, GlobalCom2 achieves superior performance-efficiency trade-offs, maintaining 90% of the original performance while dramatically reducing peak memory usage by 40% and boosting inference throughput by 1.8×1.8$\times$1.8 ×. Detailed efficiency analysis can be found in Appendix 10.4.
Method VQAT POPE MME SQA MM-Vet Avg.
Upper Bound, 2880 Tokens
Vanilla 64.9 86.5 1519.0 70.1 43.9 100.0%
Ratio=25%, Retain up to 720 Tokens
Local only 60.6 87.4 1473.3 67.6 39.6 95.6%
Global only 60.2 86.4 1488.5 67.9 37.8 94.7%
Global and Local 60.9 87.6 1493.5 68.1 40.7 96.7%
Table 3: Effects of different token evaluation metrics.
5.5 Visualizations
Figure 6 visualizes token compression on VQAT, where FasterVLM [52] and our GlobalCom2 are compared under aggressive compression ($R=10\%$). FasterVLM assumes uniform information richness across crops and preserves the same ratio of tokens for each local crop, evaluating token importance only from local perspectives. This leads to both information loss and redundancy retention, as evidenced by the redundant camera metal parts preserved in the last two crops of the first case. In contrast, our GlobalCom2 adopts a global perspective to assess actual informativeness of each crop and adaptively adjusts compression intensity accordingly. This enables our method to preserve crucial visual details while maximizing redundancy removal, achieving 1.9% higher accuracy than FasterVLM on VQAT under $R=10\%$. More cases in Appendix 8 show GlobalCom2’s consistent effectiveness in adaptive token preservation across both global and local views.
@@ 5.5 Visualizations
SparseVLM 5.4 (↓87.1%) 24.2 (↑5.2%) 5.9 (1.6$\times$) 82.7%
FasterVLM 3.8 (↓90.9%) 13.6 (↓40.1%) 6.7 (1.8$\times$) 89.8%
GlobalCom2 3.8 (↓90.9%) 13.9 (↓40.0%) 6.7 (1.8$\times$) 91.5%
Table 4: Efficiency comparisons. “Memory”: peak GPU memory; “Throughput”: POPE samples/second; “Performance”: average score on 10 multimodal benchmarks.
@@ 6 Conclusion
Token compression has achieved significant progress in accelerating LVLM inference. When applying existing methods to HR-LVLMs with dynamic tiling strategy, these methods treat global thumbnails and local crops uniformly, overlooking their inherent characteristics. Through analyzing HR-LVLMs with dynamic tiling, we reveal distinct roles between thumbnails and crops, and observe varying information densities across crops. Based on these findings, we propose GlobalCom2, a plug-and-play token compression framework that operates on a “global-to-local” guided principle, adaptively preserves informative regions as well as minimizes redundancy. Experiments show that GlobalCom2 achieves superior performance and efficiency, significantly outperforming existing baselines.
Supplementary Material
In the appendix, we provide theoretical FLOPs calculation in Section 7, more discussions about GlobalCom2 without [CLS] token in Section 8, detailed benchmarks introduction and implementation details in Section 9, more additional experiments and analysis in Section 10, and detailed explanation of our GlobalCom2 in Section 11.
7 Theoretical Complexity Analysis
GlobalCom2 compresses a large number of visual tokens for HR-LVLMs, thereby reducing their computational costs. Below, we analyze the theoretical computational complexity of HR-LVLMs in both the prefill stage and the decoding stage.
@@ 7 Theoretical Complexity Analysis
8 More Discussions about Token Evaluation without [CLS].
Some Vision Encoders like SigLIP [50] do not own [CLS] tokens, rendering existing token compression methods that rely on [CLS] tokens (e.g., FasterVLM [52]) ineffective. Therefore, we explore alternative token importance evaluation approaches for such encoders. This enables GlobalCom2 to measure information content in each crop and adaptively allocate token budgets, preserving tokens with higher information richness.
@@
$c_{i}=\cos(\mathbf{x}_{i},\mathbf{g})=\frac{\mathbf{x}_{i}\cdot\mathbf{g}}{\|\mathbf{x}_{i}\|\|\mathbf{g}\|},$ (10)
@@
Method VQAT POPE MME MM-Vet MMB SQA Avg.
@@
9 Detailed Experimental Settings
@@
  • GQA [17]: Contains 113,018 real-world images with scene graphs for structured visual reasoning.
  • VizWiz [15]: Features over 39,000 questions from blind users with lower-quality images and conversational style.
  • SQA [33]: Contains 21,000 scientific questions covering 26 topics and 379 different reasoning skills.
  • VQAT [39]: Features 28,000 images focusing on comprehension of text embedded in natural scenes.
@@
  • AI2D [19]: Contains 5,000 high-resolution scientific diagrams paired with visual-spatial reasoning questions.
  • MMStar [7]: Contains 12,000 high-resolution images for spatial, temporal, and commonsense reasoning.
@@
  • POPE [23]: Features 14,000 images with binary questions for detecting object hallucination phenomena.
  • MME [12]: Contains 15,000 images across 14 perceptual and cognitive reasoning subtasks.
  • MMB [32]: Features 20,000 multiple-choice questions designed for robust visual reasoning.
  • MMBCN [32]: Chinese counterpart of MMB with 20,000 questions for cross-lingual evaluation.
  • MM-Vet [48]: Features 16,000 questions across recognition, OCR, knowledge, and spatial reasoning.
  • SEED-Bench [20]: Contains 19,000 human-annotated questions across 12 evaluation dimensions.
Baseline Models. We select LLaVA-NeXT [29] and LLaVA-OneVision (SI) [21] as our baseline models, and follow the same inference setting as the original paper as it is publicly available [^1] LLaVA-NeXT: https://github.com/haotian-liu/LLaVA/blob/main/docs/Evaluation.md, LLaVA-OneVision: https://github.com/LLaVA-VL/LLaVA-NeXT/blob/main/docs/LLaVA_OneVision.md.. LLaVA-NeXT and LLaVA-OneVision share a common three-component architecture: a pre-trained vision encoder, a large language model (LLM) backbone, and a two-layer MLP projector bridging the two. Specifically, LLaVA-NeXT employs CLIP-ViT-L-336px [36] and Vicuna-v1.5 for vision and language modeling respectively, while LLaVA-OneVision utilizes SigLIP-So400m-Patch14-384 [50] and Qwen2 [44]. To effectively process high-resolution visual inputs, both models adopt flexible grid configurations - LLaVA-NeXT supports $\{2\times 2,1\times\{2,3,4\},\{2,3,4\}\times 1\}$ with maximum $5\times 576$ grid tokens, while LLaVA-OneVision allows $\{1\times 1,...,6\times 6\}$ grids up to $10\times 729$ tokens.
Comparison Methods. We compare our GlobalCom2 with three dominant baseline methods:
  • FastV [8] performs one-time token pruning after the selected layer of LLM based on attention map analysis.
  • SparseVLM [53] ranks token importance by text-visual attention maps, pruning through pre-selected text prompts to reduce attention noise.
  • FasterVLM [52] reranks visual tokens by [CLS] attention scores from CLIP-ViT and preserves top-$k$ tokens.
All experiments are conducted on NVIDIA A100-SXM4-80GB GPUs.
Method VQAv2 GQA VizWiz SQA VQAT POPE MME MMB MMBCN MM-Vet Average
@@
FasterVLM [52] 74.5 58.1 52.8 70.5 58.0 81.6 1386.2 61.7 53.5 38.3 88.2%
GlobalCom2 77.0 58.3 52.5 71.8 60.3 82.4 1399.5 65.0 58.5 37.9 90.3%
Table 6: Comparison with other training-free LVLM token compression methods with LLaVA-NeXT-13B across 10 benchmarks.
10 Additional Experiments
10.1 Comparisons on LLaVA-NeXT-13B.
Considering that both FasterVLM and GlobalCom2 perform token compression at the vision encoding stage, Table 6 focuses on comparing these two methods. The results reveal several key findings: (1) Our GlobalCom2 outperforms FasterVLM on most benchmarks, maintaining above 90% of uncompressed performance across different retention ratios on both LLaVA-NeXT-7B and 13B models. (2) On visual text understanding (VQAT), GlobalCom2 shows particular strength at low retention ratios, surpassing FasterVLM by 2.6% at $R=10\%$, thanks to its effective preservation of object and textual information. (3) While FasterVLM performs slightly better on general visual tasks (VQAv2, MMB, MMBCN) at high retention ratios due to its uniform token compression, GlobalCom2 demonstrates superior degradation resistance at low ratios. For example, when $R$ drops from 25% to 10%, FasterVLM’s performance significantly decreases by 5.9% and 8.6% on MMB and MMBCN, while GlobalCom2 only drops by 2.9% and 3.7%, benefiting from our carefully designed global view-guided compression strategy.
Figure 7: Hyper-parameter sensitivity analysis of $\tau$ and $\alpha$. Hyper-parameter $\tau$ controls the sharpness of probability distribution in softmax when assessing local crop informativeness, while hyper-parameter $\alpha$ determines the token retention threshold where higher values indicate stronger reliance on global guidance.
10.2 Results on LLaVA-OneVision-0.5B.
We adopt $s_{i}^{sim}$ proposed in Section 8 as the measure for crop information richness and token informativeness, and evaluate our GlobalCom2 on the LLaVA-OneVision model with more local crops across multiple benchmarks.
Our experimental results show that GlobalCom2 maintains satisfactory performance even under low retention ratios ($R=25\%,15\%,10\%$). Particularly, GlobalCom2 achieves promising results on AI2D and MMStar benchmarks which involve high-resolution images. Interestingly, we observe that on the SQA benchmark, the model performance shows minimal degradation as token retention rate decreases. We attribute this to SQA’s lower dependency on visual signals and higher reliance on LLM capabilities. Given that LLaVA-OneVision employs the powerful Qwen2 [44] as its LLM decoder, it consistently performs well on this benchmark regardless of retention ratio.
@@ 10 Additional Experiments
Table 7: Results of adopting GlobalCom2 on LLaVA-OV-0.5B.
10.3 Sensitivity Analysis of Hyper-parameters.
Hyper-parameter $\tau$ serves as the temperature in the softmax function when assessing local crop significance, controlling the “sharpness” of probability distribution. A smaller $\tau$ leads to a sharper distribution, amplifying the differences in global visual importance among local crops. As shown in the first row of Figure 7, our GlobalCom2 demonstrates robust performance across most benchmarks, particularly on VQAT [39] and POPE [23], with varying $\tau$ validating our design principle for HR-LVLMs of allocating retention ratios based on each local crop’s global importance. We empirically set $\tau=10$ to achieve optimal performance across most benchmarks.
Hyper-parameter $\alpha$ determines the criterion for token retention in local crops, where a larger $\alpha$ indicates stronger dependence on global guidance for token retention in local crops. Similar to $\tau$, the second row of Figure 7 shows that our GlobalCom2 exhibits consistent effectiveness across most benchmarks under different $\alpha$ settings, substantiating our dual-perspective token retention strategy of local crops by considering their significance at both global and local levels for HR-LVLMs.

### v5

@@ 5.4 Efficiency Analysis
Table 6 compares inference efficiency among SparseVLM, FasterVLM and GlobalCom2. Unlike SparseVLM which requires explicit attention scores in LLM and is naturally incompatible with FlashAttention (Dao et al. 2022), FasterVLM and our GlobalCom2 enable efficient computation before LLM decoding for efficient computation. As a plug-and-play solution, GlobalCom2 achieves superior performance-efficiency trade-offs, maintaining 90% of the original performance while dramatically reducing peak memory usage by 40% and boosting inference throughput by $1.8\times$. Detailed analysis is in Appendix Table 10.
5.5 Compression Visualizations
Figure 7 compares compression processes of different methods at $R=10\%$ on VQAT, revealing over-compression issues in baselines: (i) Positional Bias: FastV exhibits clear positional bias, allocating more tokens ($3\times$) to later-positioned crops regardless of visual content. (ii) Uniform Compression: FasterVLM applies uniform compression across all crops, treating them as equally important. It fails to preserve critical information in some regions while retaining redundancy in others. In contrast, GlobalCom2 globally assesses crop informativeness and adaptively adjusts compression ratios, preserving crucial details while eliminating redundancy. See Appendix Figure 11 for more examples.
@@ 5.5 Compression Visualizations
[…]
Table 6: Efficiency comparisons. “Memory”: peak GPU memory; “Throughput”: POPE samples/second; “Performance”: average score on eight multi-modal benchmarks.
@@ 6 Conclusion
Token compression has achieved significant progress in accelerating LVLM inference. When applying existing methods to HR-LVLMs with dynamic cropping, these methods treat global thumbnails and local crops uniformly, overlooking their inherent characteristics. Through analyzing HR-LVLMs with dynamic cropping, we reveal distinct roles between thumbnails and crops, and observe varying information densities across crops. Based on these findings, we propose GlobalCom2, a plug-and-play token compression framework that operates on a “global-to-local” guided principle, adaptively preserving informative regions while minimizing redundancy. Experiments show that GlobalCom2 achieves superior performance and efficiency across benchmarks, significantly outperforming existing baselines.
Appendix
In the appendix, we provide theoretical FLOPs calculation, more discussions about content-agnostic positional bias, GlobalCom2 without [CLS] token, detailed benchmarks introduction and implementation details, more additional experiments and analysis, and detailed algorithm.
Appendix A Theoretical Complexity Analysis
GlobalCom2 compresses visual tokens for HR-LVLMs, thereby reducing their computational costs. Below, we analyze the theoretical computational complexity of HR-LVLMs in both the prefill stage and the decoding stage.
@@ Appendix A Theoretical Complexity Analysis
Appendix B More Discussions about Content-agnostic Positional Bias.
Figure 9: Detailed analysis of content-agnostic positional bias. LLM attention-guided methods (e.g., FastV (Chen et al. 2024b), PDrop (Xing et al. 2025)) assign disproportionately high scores (bars) to tokens in later positions, regardless of their actual semantic content or input order (second row: sequential crop input; third row: reverse crop input).
Figure 9 reveals the inherent content-agnostic positional bias of the LLM attention-guided method FastV. Regardless of input order (sequential or reverse), FastV consistently assigns higher attention scores to later-positioned tokens when measuring visual token importance via LLM’s second-layer attention. In extreme compression settings ($R=10\%$), this bias causes severe over-compression—preserving later tokens while discarding potentially more informative earlier ones, resulting in sub-optimal performance and significant multi-modal hallucinations (14.8-point drop on POPE).
Appendix C Discussions about LVLMs without [CLS].
Some Vision Encoders like SigLIP (Zhai et al. 2023) do not own [CLS] tokens, rendering existing token compression methods that rely on [CLS] tokens (e.g., FasterVLM (Zhang et al. 2024)) ineffective. Therefore, we explore alternative token importance evaluation approaches. This enables GlobalCom2 to measure information content in each crop and adaptively allocate token budgets, preserving tokens with higher information richness.
@@
$c_{i}=\text{sim}(\mathbf{x}_{i},\mathbf{g})=\frac{\mathbf{x}_{i}\cdot\mathbf{g}}{\|\mathbf{x}_{i}\|\|\mathbf{g}\|},$ (11)
@@
Method VQAT POPE MME $M$-Vet MMB SQA Avg.
@@
Appendix D Detailed Experimental Settings
@@
  • GQA (Hudson and Manning 2019): Contains 113,018 real-world images from Visual Genome with structured scene graphs for compositional visual reasoning. Features 22M diverse questions with functional program representations, focusing on multi-hop reasoning and semantic understanding with balanced answer distributions to reduce dataset biases.
  • VizWiz (Gurari et al. 2018): Includes over 31,000 visual questions from blind users who took photos with phones and recorded spoken questions. Images are often poor quality due to lighting and framing issues, with conversational-style questions that may be unanswerable, representing real-world accessibility challenges.
  • SQA (Lu et al. 2022): Contains 21,208 multimodal multiple-choice scientific questions from elementary and high school curricula, covering natural science, language science, and social science. Features 26 topics, 127 categories, and 379 different reasoning skills with detailed lectures (83.9%) and explanations (90.5%) for answers.
  • VQAT (Singh et al. 2019): Features 28,408 high-resolution images from OpenImages with 45,336 questions focusing on reading and reasoning about text embedded in natural scenes. Requires OCR capabilities and text-based visual reasoning, with questions averaging 7.18 words in length.
@@
  • AI2D (Kembhavi et al. 2016): Contains 5,000 high-resolution scientific diagrams from textbooks paired with 16,000 multiple-choice questions requiring visual-spatial reasoning about complex scientific concepts, processes, and relationships depicted in educational illustrations.
  • MMStar (Chen et al. 2024a): Contains 1,500 challenging samples across 6 core capabilities (coarse perception, fine-grained perception, instance interaction, logical reasoning, science & technology, mathematics) designed to evaluate vision-indispensable reasoning without relying on shortcuts.
@@
  • POPE (Li et al. 2023): Features 3,000 images with 9,000 binary yes/no questions specifically designed for detecting object hallucination phenomena in large vision-language models. Uses systematic adversarial evaluation to probe model reliability and factual accuracy.
  • MME (Fu et al. 2023): Contains 2,374 high-resolution images across 14 perceptual and cognitive reasoning subtasks, including existence, count, position, color, OCR, commonsense reasoning, numerical calculation, text translation, and code reasoning. Uses yes/no questions for comprehensive evaluation.
  • MMB (Liu et al. 2024d): Features 2,974 multiple-choice questions designed for robust visual reasoning evaluation across 20 ability dimensions including object localization, attribute recognition, scene understanding, and spatial relationship reasoning.
  • MMB-CN (Liu et al. 2024d): Chinese counterpart of MMBench with 2,974 questions translated and culturally adapted for cross-lingual evaluation, maintaining the same 20 ability dimensions while incorporating Chinese cultural contexts and linguistic nuances.
  • MM-Vet (Yu et al. 2024): Features 218 high-quality questions across 6 core VL capabilities (recognition, OCR, knowledge, language generation, spatial awareness, math) with complicated multi-modal reasoning chains requiring integration of multiple skills for comprehensive evaluation.
  • SEED-Bench (Li et al. 2024a): Contains 19,242 human-annotated multiple-choice questions across 12 evaluation dimensions covering both image and video understanding tasks. Features hierarchical evaluation from basic perception to complex reasoning with balanced difficulty distribution.
  • MVBench (Li et al. 2024b): Defines 20 video understanding tasks requiring deep temporal comprehension beyond single-frame analysis, including dynamic scene understanding, temporal action localization, multi-object tracking, and causal reasoning across video sequences.
  • LongVideoBench (Wu et al. 2024): Focuses on long-context video understanding with 3,763 videos up to one hour duration and 6,678 questions across 17 categories including plot understanding, character analysis, temporal reasoning, and comprehensive video summarization.
  • MLVU (Zhou et al. 2025): Features videos from 3 minutes to 2 hours with 9 tasks including topic reasoning, video summarization, needle-in-a-haystack retrieval, and ego-centric analysis, designed to evaluate long-form video comprehension and temporal memory capabilities.
  • VideoMME (Fu et al. 2024): Comprises 900 videos and 2,700 questions across six domains (Knowledge, Film & Television, Sports, Life, Subtitles, Games) with durations from 11 seconds to 1 hour, featuring both with-subtitle and without-subtitle evaluation settings.
Method VQAv2 GQA SQA VQAT POPE MME MMB MMB-CN Average
@@
FasterVLM(2024.12) 74.5 58.1 70.5 58.0 81.6 1386.2 61.7 53.5 88.6%
GlobalCom2 77.0 58.3 71.8 60.3 82.4 1399.5 65.0 58.5 90.9%
Table 8: Comparison with FasterVLM on LLaVA-NeXT-13B across multiple benchmarks.
Baseline Models. We select LLaVA-NeXT (Liu et al. 2024b) and LLaVA-OneVision (SI) (Li et al. 2025) as our HR-LVLMs models, and follow the same inference setting as the original paper as it is publicly available [^2] LLaVA-NeXT: https://github.com/haotian-liu/LLaVA/blob/main/docs/Evaluation.md, LLaVA-OneVision: https://github.com/LLaVA-VL/LLaVA-NeXT/blob/main/docs/LLaVA˙OneVision.md.. LLaVA-NeXT and LLaVA-OneVision share a common three-component architecture: a pre-trained vision encoder, a large language model (LLM) backbone, and a two-layer MLP projector bridging the two. Specifically, LLaVA-NeXT employs CLIP-ViT-L-336px (Radford et al. 2021) and Vicuna-v1.5 for vision and language modeling respectively, while LLaVA-OneVision utilizes SigLIP-So400m-Patch14-384 (Zhai et al. 2023) and Qwen2 (Yang et al. 2024). To effectively process high-resolution visual inputs, both models adopt flexible grid configurations - LLaVA-NeXT supports $\{2\times 2,1\times\{2,3,4\},\{2,3,4\}\times 1\}$ with maximum $5\times 576$ grid tokens, while LLaVA-OneVision allows $\{1\times 1,...,6\times 6\}$ grids up to $10\times 729$ tokens. For hyper-parameters, $\tau$ and $\alpha$ are respectively set as 10 and 0.5 for benchmark evaluations.
We also evaluate two VideoLLMs: LLaVA-OneVision and Qwen2-VL (Wang et al. 2024). LLaVA-OneVision unifies image and video tasks by encoding videos as long image-style token sequences, enabling strong zero-shot video understanding. Qwen2-VL adopts Naive Dynamic Resolution and Multimodal Rotary Position Embedding to achieve long video understanding (>>20 min).
Comparison Methods. We compare our GlobalCom2 with below dominant LVLM token compression methods:
  • FastV (Chen et al. 2024b) performs one-time pruning after the selected LLM layer based on attention weights between vision tokens and the last token.
  • PDrop (Xing et al. 2025) introduces progressive token dropping using similar token selection metrics as FastV, forming a pyramid-like token structure that balances efficiency and performance.
  • SparseVLM (Zhang et al. 2025) ranks token importance by text-visual attention maps, pruning through pre-selected text prompts to reduce attention noise.
  • PruMerge (Shang et al. 2025) integrates token pruning and merging by removing less important tokens using [CLS] attention weights with patch tokens and clustering retained tokens based on key similarity.
  • FasterVLM (Zhang et al. 2024) re-ranks visual tokens by using [CLS] attention scores with all patch tokens from ViT and preserves top-$k$ tokens.
We also compare with two VideoLLM-specific methods:
  • DyCoke (Tao et al. 2025) groups video frames using a four-frame sliding window and performs temporal token merging within each window.
  • FrameFusion (Fu et al. 2025) adopts a two-stage strategy: first merging tokens across frames based on visual similarity, then selecting tokens within each frame based on importance.
All experiments in this work are conducted on NVIDIA A100-SXM4-80GB GPUs.
Appendix E Additional Experiments
@@
Table 9: Results of GlobalCom2 on LLaVA-OV-0.5B.
E.1 Comparisons on LLaVA-NeXT-13B
Considering that both FasterVLM and GlobalCom2 perform token compression at the vision encoding stage, Table 8 focuses on comparing these two methods on larger HR-LVLM LLaVA-NeXT-13B. The results reveal several key findings: (i) Our GlobalCom2 outperforms FasterVLM on most benchmarks, maintaining above 90% of uncompressed performance across different retention ratios on both LLaVA-NeXT-7B and 13B models. (ii) On visual text understanding (VQAT), GlobalCom2 shows particular strength at low retention ratios, surpassing FasterVLM by 2.3% at $R=10\%$, thanks to its effective preservation of object and textual information. (iii) While FasterVLM performs slightly better on general visual tasks (VQAv2, MMB, MMB-CN) at high retention ratios due to its uniform token compression, GlobalCom2 demonstrates superior degradation resistance at low ratios. For example, when $R$ drops from 25% to 10%, FasterVLM’s performance significantly decreases by 5.9% and 8.6% on MMB and MMB-CN, while GlobalCom2 only drops by 2.9% and 3.7%, benefiting from our carefully designed “global-to-local” guided compression strategy.
E.2 Results on LLaVA-OneVision-0.5B
We adopt $s_{i}^{sim}$ proposed in Section C as the measure for crop information richness and token informativeness, and evaluate our GlobalCom2 on the LLaVA-OneVision model with more local crops across multiple benchmarks.
Results show that GlobalCom2 maintains satisfactory performance even under low retention ratios ($R=25\%,15\%,10\%$). Particularly, GlobalCom2 achieves promising results on AI2D and MMStar benchmarks which involve high-resolution images. Interestingly, we observe that on the SQA benchmark, the model performance shows minimal degradation as token retention rate decreases. We attribute this to SQA’s lower dependency on visual signals and higher reliance on LLM capabilities. Given that LLaVA-OneVision employs the powerful Qwen2 (Yang et al. 2024) as its LLM decoder, it consistently performs well on this benchmark regardless of retention ratio.
Figure 10: Hyper-parameter sensitivity analysis of $\tau$ and $\alpha$. Hyper-parameter $\tau$ controls the sharpness of probability distribution in softmax when assessing local crop informativeness. Hyper-parameter $\alpha$ determines the token retention threshold where higher values indicate stronger reliance on global guidance.